40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 5 12:27:27 2015, manasa, Update, General, FOL fiber box on the PSL table 
    Reply  Fri Mar 6 14:29:59 2015, manasa, Update, General, FOL troubleshooting 
       Reply  Mon Mar 9 14:43:31 2015, manasa, Update, General, FOL troubleshooting FOL_trouble.pdfIMAG0341.jpgIMAG0342.jpg
          Reply  Tue Mar 10 17:48:07 2015, manasa, Update, General, FOL troubleshooting 
             Reply  Wed Mar 11 19:48:25 2015, Koji, Update, General, FOL troubleshooting 
                Reply  Thu Mar 12 11:57:38 2015, Koji, Update, General, FOL troubleshooting 
Message ID: 11137     Entry time: Thu Mar 12 11:57:38 2015     In reply to: 11135
Author: Koji 
Type: Update 
Category: General 
Subject: FOL troubleshooting 

BTW, during this trouble shoot, we looked at the IR beatnote spectrum between the Xend and the PSL.
It showed a set of sidebands at ~200kHz, which is the modulation frequency.
There was another eminent component present at ~30kHz.
I'm afraid that there is some feature like large servo bump, a mechanical resonance, or something else, at 30kHz.

We should check it. Probably it is my job.

ELOG V3.1.3-